OpenStudy (anonymous):

Identify the missing symbols: 2/5 ? 3/5 ? 9/15 > and = < and = < and < < and >

OpenStudy (anonymous):

what is this ? a pre-algebra ?

OpenStudy (anonymous):

Actually, it is too easy :( I don't know how to explain. OK, let do it as: Firstly. make all fractions has the same denominator, in this case is 5. \[\frac{ 2 }{ 5 } ; \frac{ 3 }{ 5 } ; and ; \frac{ 9 }{ 15 } = \frac{ 3*3 }{ 3*5 } = \frac{ 3 }{ 5 }\] For 3 fractions with same denominator, you just compare to numerator, so the answer is \[\frac{ 2 }{ 5 } < \frac{ 3 }{ 5 } = \frac{ 3 }{ 5 } or \frac{ 2 }{ 5 } < \frac{ 3 }{ 5 } = \frac{ 9 }{ 15 }\]

OpenStudy (phi):

If you cut a pie into 5 equal pieces 2/5 means 2 of the 5 pieces 3/5 means 3 of the 5 pieces which is more 2/5 or 3/5 ?